a plane traveling with the wind can travel 1170 miles in 3 hours. Against the same wind the plane can travel 1550 miles in 5 hours. How fast are the plane and the wind traveling?>
Found 3 solutions by TimothyLamb, Alan3354, Danielshinmath:
Answer by TimothyLamb(4379)

x = speed of airplane in still air
y = speed of wind
---
s = d/t
---
downwind:
x + y = 1170/3
x + y = 390
---
upwind:
x - y = 1550/5
x - y = 310
---
put the system of linear equations into standard form
---
x + y = 390
x - y = 310
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = speed of airplane in still air = 350 mph
y = speed of wind = 40 mph
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by Alan3354(69443)

Find the 2 groundspeeds, upwind and downwind.
The plane's airspeed is the average of the 2.
----
windspeed is the difference between airspeed and groundspeed.
==================
No reason to make it a doctoral dissertation.

Answer by Danielshinmath(31)

Find the mph for both trips. Make the plane x and the wind y. So if the plane travels with the wind, it goes 390 mph. x+y=390. If the plane travels against the wind, it goes 310 mph. 310=x-y.
do x+y=390
- x-y=310
2y= 80, so y or the wind speed is 40mph. and the plane speed is 350 mph.